OpenStudy (anonymous):

If F(x)= the integral from (0,x) the squareroot of (t^3+1)dt then F' (2)

OpenStudy (anonymous):

\[\int\limits_{0}^{x}\sqrt{t ^{3}+1}\]

OpenStudy (amistre64):

might be better like this ... \[F(t)=\int F'(t)dt\] \[\int_{0}^{x}F'(t)dt=F(x)-F(0)\] \[\frac{d}{dx}F(x)-F(0)=F'(x)x' - F'(0)0'\]

OpenStudy (amistre64):

so my guess is; sqrt(2^3 +1) = sqrt(9) = 3

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yw